OpenStudy (anonymous):

]In a school of 200 students, 80 students are in the band, 160 students are on sports teams, and 60 students participate in both activities. How many students in the school are neither in band nor on a sports team?

OpenStudy (anonymous):

I said 20..is that correct

OpenStudy (aum):

correct. What method did you use?

OpenStudy (aum):



OpenStudy (aum):

Students neither in band nor on a sports team is: 200 - (20 + 60 + 100) = 200 - 180 = 20